# De-Hallucinator: Iterative Grounding for LLM-Based Code Completion

### v1

@@ 3.2. Overview and Main Algorithm
Figure 4 gives a high-level overview of the approach, and Fig. 5 shows each step of the approach on our running example from Fig. 1. Using the code before the cursor location as an initial prompt, De-Hallucinator queries the model to get an initial completion. Because the model is not aware of project-specific APIs, the completion is likely to not make best use of these APIs. For the example in Fig. 5, the initial completion refers to a non-existing API doc.relevance.
To help the model predict a better completion, De-Hallucinator retrieves a ranked list of related API references in two ways. The first is to retrieve API references similar to the initial prompt. The second approach is to retrieve them based on similarity to the preceding code and the initial completion of the model. In Fig. 5, this step retrieves a reference to the relevance function defined in utils.py. Note that in this example the retrieved API reference is the same when retrieved based on the initial prompt and based on the initial completion. For efficient retrieval of APIs, De-Hallucinator analyzes the project in advance and indexes all APIs in the project. The approach then uses the related API references to construct an augmented prompt, and queries the model again. In the example, the second query returns the correct completion, as shown at the bottom of Fig. 5. De-Hallucinator continues the retrieval based on the previous completion by the model until either reaching a fixed point, i.e., the model generates the same completion twice in a row, or until exhausting a configurable maximum number of queries, $k$.
%**** main.tex Line 300 ****…
def search(ds: DataStore, keyword: str, top_k: int) -> List[str]:
"""Returns the top_k most relevant documents that contain the keyword sorted by relevance."""
docs = ds.find_by_keyword(keyword)
def search(ds: DataStore, keyword: str, top_k: int) -> List[str]:
"""Returns the top_k most relevant documents that contain the keyword sorted by relevance."""
docs = ds.find_by_keyword(keyword)
[…]
relevance(document: str, keyword: str) -> float # Returns the relevance of the document to the keyword.
# API Reference:
# relevance(document: str, keyword: str) -> float # Returns the relevance of the document to the keyword.
def search(ds: DataStore, keyword: str, top_k: int) -> List[str]:
"""Returns the top_k most relevant documents that contain the keyword sorted by relevance."""
docs = ds.find_by_keyword(keyword)
def search(ds: DataStore, keyword: str, top_k: int) -> List[str]:
"""Returns the top_k most relevant documents that contain the keyword sorted by relevance."""
docs = ds.find_by_keyword(keyword)
[…]
@@
Data: Incomplete code $c$, model $m$, budget $k$
Result: List $\mathbb{C}$ of completions
[…]
$\mathbb{C}$ = $[c^{\prime}]$;
[…]
$\mathbb{C}=\mathbb{C}+[c^{\prime}]$;
for $i\in\{1,...,k-1\}$ do
[…]
       $\mathbb{C}=\mathbb{C}+[c^{\prime}]$;
[…]
             break;
       end if
[…]
end for
return $\mathbb{C}$
Algorithm 1 Main algorithm of De-Hallucinator.
Algorithm 1 shows the pseudo-code of the main algorithm of De-Hallucinator. The call to retrieveRelevantAPIrefs corresponds to the “Retrieval of related APIs” component in Fig. 4, and the call to constructPrompt matches the “Prompt construction” component in the figure. The following presents these two components, as well as the static pre-analysis that enables an efficient retrieval of related APIs, in detail.
3.3. Static Pre-Analysis
To ensure that the retrieval of API references does not unnecessarily slow down the code completion, De-Hallucinator has a preprocessing phase that indexes the current project for fast retrieval. We use API reference throughout this paper to refer to a piece of text extracted from the project’s code, which can be added to the prompt to provide further information about a project-specific API.
@@ 3.3. Static Pre-Analysis
  • An attribute reference, which is the qualified name of a self attribute assigned to in the __init__ function.
@@
Source
API reference
Type of API reference
DataStore.py in Fig. 1
DataStore.find_by_keyword(self, keyword: str) -> List[str] # Returns all documents that contain the keyword.
Function reference
utils.py in Fig. 1
relevance(document: str, keyword: str) -> float # Returns the relevance of the document to the keyword.
Function reference
DataStore.py in Fig. 1
class DataStore() # Data structure for storing documents.
Class reference
DataStore.py in Fig. 1
DataStore.documents
Attribute reference
Table 1 shows some of the API references extracted from our example project.
To extract API references from a project, the preprocessing phase of De-Hallucinator builds on CodeQL. [^5] https://codeql.github.com/ CodeQL is a query language and framework for statically analyzing source code, which provides a declarative approach for querying a code base. The framework uses static analysis to create a database and run the queries. Using CodeQL allows adding support for other languages with minimal effort. We implement two CodeQL queries, one for function references, and the other for class references and attribute references. Alternatively to our approach for gathering API references, an IDE-based implementation could reuse information about the current project that is computed by the static code indexing performed by an IDE anyway.
Because an effective code completion technique must provide suggestions quickly, De-Hallucinator indexes the extracted API references for fast retrieval, which takes two steps. The first step of the indexing is to embed API references into a vector space. Formally, we need an embedding function, $E$, for which $E(c)=v_{c}\in\mathbb{R}^{d}$, such that, for two code pieces $c_{1}$ and $c_{2}$, the cosine similarity of their embeddings, $v_{c_{1}}\cdot v_{c_{2}}/(|v_{c_{1}}||v_{c_{2}}|)$, approximates the semantic similarity of $c_{1}$ and $c_{2}$. To this end, we embed the API references into a vector space using Sentence-BERT (Reimers and Gurevych, 2019), a BERT-based model designed for measuring the semantic similarity between sentences. We use a variant of this model that is pre-trained on code. [^6] https://huggingface.co/flax-sentence-embeddings/st-codesearch-
       distilroberta-base
The model maps sentences, or in our case lines of code, into a dense vector representation of size 768. Because our approach uses the embedding function as a black-box, any other embedding model or similarity-preserving vector representation (Wainakh et al., 2021) could also be used instead, e.g., GloVe (Pennington et al., 2014), BERT (Devlin et al., 2017), or FastText (Bojanowski et al., 2017). The second step, after embedding the API references, is to index the normalized vectors ($v_{c}/|v_{c}|$ for all $c\in\text{API references}$) in a Ball Tree. [^7] https://scikit-learn.org/stable/modules/neighbors.html#ball-tree This index allows for fast retrieval of nearest neighbors.
@@ 3.4. Retrieval of Related APIs
Getting back to our running example, consider the third section in Fig. 5. It shows the API reference from our example project that the retrieval component finds to be the top-most relevant ($n=1$) for both the incomplete code (initial prompt) and the initial prediction of the LLM. Our implementation considers multiple relevant API references, i.e., $n>1$, which we omit in the example for brevity.
@@ 3.5. Prompt Construction
Given the code-to-complete and a list $R$ of API references that may enable the LLM to accurately complete the code, De-Hallucinator constructs an augmented prompt for querying the model. The prompt is designed in a way that resembles “normal” code, i.e., the kind of data that the LLM has been trained on. The prompt structure we use is shown in Fig. 6. The API references are shown as a block of commented lines at the beginning of the prompt. These lines start with API Reference:, and the following lines contain the relevant API references in decreasing order of similarity to the lines in the input of the retrieval module. This commented block is followed by the original prompt clipped to fit in the prompt size.
# API Reference:
# <most relevant API reference>
# <2nd most relevant API reference>
# <|\$n\$|th most relevant API reference>
<clipped original prefix code>
Figure 6. Template for the augmented prompt.
For our running example, the fourth section in Fig. 5 shows the prompt for function search in our example, augmented with the API reference. Given the augmented prompt, the same LLM that predicted the code in Fig. 2, completes this function correctly in the last section of Fig. 5.
@@ 3.6. Integration with the LLM
De-Hallucinator is designed with minimal assumptions about the underlying LLM. The approach considers the model to be a black box that we query with a string that contains incomplete code and that returns a string with a suggested completion of the code. After receiving the completion, De-Hallucinator post-processes it to make the completion syntactically correct and to remove any extra completions that are out of the scope of the current completion. For example, during the completion of a function body, when the model generates a long completion, it is possible that the completion closes the scope of the current function, and starts defining a new function, or write code outside of the function body. Since our approach is focused towards API usages, the post-processing aims to mitigate these issues by cutting off the generated completion at the end of the line that completes the API usage.
@@ 4. Implementation
We implement the De-Hallucinator approach into a Python-based tool. For statically extracting API references from the current project, the implementation builds on CodeQL, and because we target Python code, on the Python language support of CodeQL, which offers easy access to the classes, functions, etc. in a code base. To access the LLMs, we use the HuggingFace transformers library, [^8] https://huggingface.co/transformers/ making our implementation compatible with many publicly available models. Adapting our implementation to other models requires only to adjust the prompt size of the model and to select other parameters passed to its API.
@@ 1. RQ1:
    How much does De-Hallucinator improve code completions compared to querying the LLM with the default prompt only?
@@ 3. RQ3:
    How do the hyperparameters of De-Hallucinator affect the completions?
@@ 5.1.1. LLMs and Baseline
We evaluate on four state-of-the-art LLMs: CodeGen (Nijkamp et al., 2022) with 2.7B parameters (Salesforce/codegen-2B-mono), CodeGen 2.5 (Nijkamp et al., 2023) with 7B parameters (Salesforce/codegen25-7b-mono), UniXCoder (Guo et al., 2022a) with 125M parameters (microsoft/unixcoder-base), and StarCoder+ (Li et al., 2023b) with 15.5B parameters (bigcode/starcoderplus). The reasons for selecting these models are (i) that, unlike the models offered by OpenAI, they are freely available and hence allows for building upon our work in the future, and (ii) they cover a variety of parameter sizes, model architectures, and pre-training processes. We leave all parameters of the models at their defaults, except for the maximum new tokens parameter, which we set to 256 to allow for longer completions. As a baseline, we query the models with a prompt that contains all the code preceding the cursor. In case this prompt exceeds the maximum prompt size of 2,048 tokens, we truncate the prompt from the beginning.
5.1.2. Dataset and Ground Truth
Table 2. List of Python projects used for the evaluation.
Project (owner/name) Commit
Description
LoC $*{}^{*}$
graphql-python/graphene 57cbef6
GraphQL wrapper
9,484 7.7k
geopy/geopy ef48a8c
Client for geocoding web services
10,000 3.9k
nvbn/thefuck ceeaeab
CLI tool to correct commands
12,181 77.1k
${}^{**}$ ba1ed7a
Websocket server & client library
14,186 4.5k
arrow-py/arrow 74a759b
Library for easier use of date & time
14,402 8.3k
lektor/lektor be3c8cb
Static CMS framework
16,852 3.7k
Parsely/streamparse aabd9d0
Python interface for job scheduling and distribution
26,214 1.5k
Supervisor/supervisor ca54549
Process control for Unix-like systems
29,860 7.8k
mwaskom/seaborn f9827a3
Statistical data visualization library
37,367 10.7k
psf/black ef6e079
Python code formatter
106,005 32.2k
scikit-learn/scikit-learn f3c6fd6
Machine learning framework for Python
193,863 54.0k
$*{}^{*}$ As of May 1, 2023
${}^{**}$ The project has been moved to python-websockets/websockets
With the goal of having a diverse set of projects in terms of size, domain, and popularity, we gather a dataset of eleven public Python projects from GitHub, shown in Table 2. These projects cover different domains, such as data visualization, machine learning, process management, job distribution, and geocoding. Likewise, the dataset has a wide range of project sizes and levels of popularity, as measured by GitHub stars.

### v3

@@ 3.1. Overview
3.1.1. Main Algorithm
Figure 3 gives an overview of the approach, which we use to illustrate the main algorithm. The top of the figure shows the traditional code generation process, where an LLM receives a prompt and generates code. We call this prompt the initial prompt. Because the model may not be aware of project-specific APIs, the output is likely to refer to some hallucinated APIs.
To help the model predict better code, De-Hallucinator refines the initial prompt using two techniques. Both of them add API references to the prompt, but they do so in different ways. First, as shown by the dotted lines, De-Hallucinator uses the initial prompt to retrieve related API references from the project. This approach is similar to retrieval-augmented generation (RAG) (Lewis et al., 2020), and we refer to the resulting prompt as the RAG prompt.
Second, as shown by the dashed lines, De-Hallucinator uses the output of the model to retrieve related API references. This technique is unique to our approach, and it is based on the observation that the model’s output often resembles the desired code, but fails to refer to the correct APIs. Retrieving suitable API references based on the model’s output can be done multiple times, and hence, we refer to the resulting prompt as the iterative prompt. In general, De-Hallucinator repeats the iterative prompt refinement, i.e., the dashed loop in the figure, until exhausting a configurable maximum number $k$ of queries to the model.
For efficient retrieval of APIs, De-Hallucinator analyzes the project in advance, as shown in the “Pre-analysis” component, and indexes all APIs of the project.
3.1.2. Example
Fig. 4 shows each step of the approach on our running example from Fig. 1. Given the initial prompt, the initial completion by the model refers to a non-existing API x.score. Then, for the RAG prompt the initial prompt is used for retrieval. In this step, as shown in Fig. 4, the reference to an already used function, find_by_keyword is retrieved. Consequently, the completion uses the wrong API, as the model still does not have knowledge of the relevance function. Next, using the initial completion by the model, De-Hallucinator retrieves a reference to the relevance function defined in utils.py. In the example, the iterative prompt results in the correct completion, as shown at the bottom of Fig. 4.
{minted}
[label=Initial prompt]python … def search(ds: DataStore, keyword: str, top_k: int) -> List[str]: docs = ds.find_by_keyword(keyword) {minted}[label=Initial completion,highlightlines=4]python def search(ds: DataStore, keyword: str, top_k: int) -> List[str]: docs = ds.find_by_keyword(keyword) return sorted(docs, key=lambda x:|x.score,|, reverse=True)[:top_k] {minted}[label=Most relevant API reference]text DataStore.find_by_keyword(self, keyword: str) -> List[str] {minted}[label=RAG prompt]python # API Reference: # DataStore.find_by_keyword(self, keyword: str) -> List[str] def search(ds: DataStore, keyword: str, top_k: int) -> List[str]: docs = ds.find_by_keyword(keyword) {minted}[label=RAG prompt completion,highlightlines=4]python def search(ds: DataStore, keyword: str, top_k: int) -> List[str]: docs = ds.find_by_keyword(keyword) return sorted(docs, key=lambda x:|x.score,|, reverse=True)[:top_k] {minted}[label=Most relevant API reference]text relevance(document: str, keyword: str) -> float {minted}[label=Iterative prompt]python # API Reference: # relevance(document: str, keyword: str) -> float def search(ds: DataStore, keyword: str, top_k: int) -> List[str]: docs = ds.find_by_keyword(keyword) {minted}[label=Iterative prompt completion,highlightcolor=bggreen,highlightlines=4]python def search(ds: DataStore, keyword: str, top_k: int) -> List[str]: docs = ds.find_by_keyword(keyword) return sorted(docs, key=lambda doc: relevance(keyword, doc), reverse=True)[:top_k] # <- Correct
@@ 3.1.2. Example
The following presents each component of De-Hallucinator in more detail, as well as how we instantiate the components for our two target tasks, code completion and test generation.
3.2. Pre-Analysis
3.2.1. General Idea
To ensure that the retrieval of API references does not unnecessarily slow down the code generation, De-Hallucinator has a preprocessing phase that indexes the current project for fast retrieval. We use API reference throughout this paper to refer to a piece of text extracted from the project’s code, which can be added to the prompt to provide further information about a project-specific API.
@@ 3.2.1. General Idea
  • An attribute reference, which is the qualified name of a self attribute assigned to in the constructor.
3.2.2. Example
Table 1 shows some of the API references extracted from our example project. Note that the API references resemble real code to maintain compatibility with any model that is trained on source code.
@@ 3.2.2. Example
Source API reference Type of API reference
DataStore.py DataStore.find_by_keyword(self, keyword: str) -> List[str] Function reference
utils.py relevance(document: str, keyword: str) -> float Function reference
DataStore.py class DataStore() Class reference
DataStore.py DataStore.documents Attribute reference
3.2.3. Application to Code Completion
For the code completion task, we use CodeQL [^5] https://codeql.github.com/ to statically extract APIs from the project source. The Python language support of CodeQL offers easy access to the classes, functions, etc. in a code base. Another benefit of CodeQL is that we can utilize databases created by GitHub for open-source projects.
3.2.4. Application to Test Generation
Since the APIs in JavaScript are in many cases only available once a module is instantiated, De-Hallucinator dynamically loads the modules and traverses them to extract API references.
Alternatively to our approaches for gathering API references, an IDE-based implementation could reuse information about the current project that is anyway computed by the static code indexing in an IDE.
@@ 3.3. Retrieval of Related APIs
3.3.1. General Idea
The retrieval module takes an input code piece and returns a ranked list of project-specific API references that are most similar to the input. To enable similarity-based search, De-Hallucinator embeds the extracted API references into a vector space. Formally, we need an embedding function, $E$, for which $E(c)=v_{c}\in\mathbb{R}^{d}$, such that, for two code pieces $c_{1}$ and $c_{2}$, the cosine similarity of their embeddings, $v_{c_{1}}\cdot v_{c_{2}}/(|v_{c_{1}}||v_{c_{2}}|)$, approximates the semantic similarity of $c_{1}$ and $c_{2}$. Recently, many models have been trained for this task (Cod, [n. d.]). Because our approach uses the embedding function as a black-box, any embedding model or similarity-preserving vector representation (Wainakh et al., 2021) can be used with De-Hallucinator, e.g., GloVe (Pennington et al., 2014), BERT (Devlin et al., 2017), or FastText (Bojanowski et al., 2017).
Given the embeddings of the API references, De-Hallucinator retrieves API references that are most similar to the input code piece. To this end, the approach embeds the input code piece and searches for the $n$ API references that minimize the cosine distance to the input code piece. The parameter $n$ specifies the number of API references to include in the prompt.
3.3.2. Example
Getting back to our running example, consider the third section in Fig. 4. It shows the API reference from our example project that the retrieval component finds to be the top-most relevant to the incomplete code (initial prompt). Even though our implementation retrieves $n$ relevant API references, we show only one in the example for brevity.
3.3.3. Application to Code Completion
For code completion, we split the given input code piece into lines, then retrieve the most relevant APIs for each line, and finally merge the $n$ most similar API references into single sorted list. The reason for retrieving the API references similar to full lines of code, as opposed to only API usages, is that we want the completion to avoid re-implementing existing code. Therefore, if there exists some API similar to a line of code that does not use any APIs, we want the approach to be able to retrieve suitable API references to generate the correct completion. We embed the API references into a vector space using Sentence-BERT (Reimers and Gurevych, 2019), a BERT-based model designed for measuring the semantic similarity between sentences. We use a variant of this model that is pre-trained on code. [^6] https://huggingface.co/flax-sentence-embeddings/st-codesearch-distilroberta-base The model maps sentences, or in our case lines of code, into a dense vector representation of size 768.
@@ 3.3.3. Application to Code Completion
3.3.4. Application to Test Generation
For test generation, we extract all API usages in a previously generated test using regular expression matching and embed them. Next, we retrieve the most relevant APIs for every API usage and return a single list. Here we retrieve API references based on API usages, and not based on lines. The reason is that as there are no ground truths for the test generation task, and the main goal is to generate passing tests. Therefore, fixing a wrong API usage is more important than avoiding re-implementation of existing code. The embedding used for indexing the API references and during retrieval is a BERT-based model available on HuggingFace. [^8] https://huggingface.co/jinaai/jina-embeddings-v2-base-code The reason for using a different model than the one used for code completion is that not all models on HuggingFace are compatible with TypeScript, so we choose a compatible model. Using mean pooling, we embed the code pieces into a vector of size 768. We use the same model for both indexing and retrieval, where during indexing each API signature is embedded and stored in a list, and during retrieval the API usage is embedded. During retrieval, we perform a linear search for the most relevant API references. Since the size of JavaScript projects are much smaller than the Python projects, there is not much benefit in using the Ball Tree data structure.
@@ 3.4. Prompt Construction
3.4.1. General Idea
Given the input and a list of at most $n$ API references that may enable the LLM to accurately generate code, De-Hallucinator constructs an augmented prompt for querying the model. The prompt is designed in a way that resembles “normal” code, i.e., the kind of data that the LLM has been trained on. The API references are augmented as a block of commented lines to the prompt. These lines start with API Reference:, and the following lines contain the relevant API references in decreasing order of similarity to the lines in the input of the retrieval module.
3.4.2. Example
For our running example, the “Iterative prompt" section in Fig. 4 shows the prompt for function search in our example, augmented with the API reference. Given the iterative prompt, the same LLM that predicted the code in Fig. 2, completes this function correctly in the last section of Fig. 4.
3.4.3. Application to Code Completion
In the code completion task, we prepend the prompt with the API references, as it minimally disrupts the structure of the existing code. See our running example for illustration.
3.4.4. Application to Test Generation
On the other hand, for test generation, as TestPilot already includes additional information into the prompt, we append the API references to the end of the additional context section. Figure 6 shows how the prompt is augmented with the API references in JavaScript, and the model’s success in generating a correct test based on that prompt.
@@ 3.5. Integration with the LLM
De-Hallucinator is designed with minimal assumptions about the underlying LLM, and the tool that solves the code generation task. The approach considers the LLM to be a black box that we query with a string, which then returns one or multiple strings with suggested code pieces. Hence, we do not fine-tune the LLM, or train a model to preform retrievals, which makes De-Hallucinator applicable to more scenarios. To that end, the instantiation of De-Hallucinator in both code completion and test generation tasks are compatible with any LLM, and our experiments with five LLMs show the flexibility of the approach.
@@ 4. Implementation
The general ideas behind De-Hallucinator are language-agnostic, and the approach can be applied to different programming languages and to different code prediction tasks. We present two implementations, one for code completion in Python and one for test generation in JavaScript. The code completion implementation is in Python and builds on top of the HuggingFace transformers library. Adapting our implementation to other models requires only to adjust the prompt size of the model and to select other parameters passed to its API. The test generation implementation is in TypeScript and builds on top of the state-of-the-art LLM-based test generator TestPilot (Schäfer et al., 2024). Like TestPilot, we use GPT-3.5-turbo as the LLM. TestPilot generates tests by going through the functions in the package under test, and for each function tries to generate a test with a simple input, which consists of a test header and the signature of the function under test. Then, a set of “prompt refiners” modify the prompt to include usage snippets, error messages, and the body of the function under test. We implement two new prompt refiners, one for the RAG prompts, and one for iterative prompts. These two refiners are only activated when a test fails with an error that is likely caused by a hallucinated API, which we detect by checking if the error message contains “is not a function” or “of undefined”.
@@ 1. RQ1:
    How much does De-Hallucinator improve the generated code compared to the baseline approachs?
@@ 3. RQ3:
    How do the hyperparameters of De-Hallucinator affect the results?
@@ 5.1. Experimental Setup
5.1.1. Tasks
Our evaluation targets two tasks, code completion and test generation. For code completion, we define the task as completing an incomplete function at the beginning of a line with an API usage, given the preceding code and the existing code in the project. This problem definition matches the common scenario of a developer implementing a function in an existing project, where the code to be written should use a project-specific API. For example, suppose that the cursor in Fig. 1 is at the beginning of the code marked with gray background. Everything above the cursor is our incomplete code $c$, and the problem is to predict the marked code $c^{\prime}$, which refers, e.g., to the project-specific relevance API. For test generation, the task is to generate tests for a given JavaScript package. This problem setting has been well established by previous work (Arteca et al., 2022; Schäfer et al., 2024).
@@ 5.1.2. LLMs and Baseline
We evaluate the code completion task on four state-of-the-art LLMs: CodeGen (Nijkamp et al., 2022) with 2.7B parameters (Salesforce/codegen-2B-mono), CodeGen 2.5 (Nijkamp et al., 2023) with 7B parameters (Salesforce/codegen25-7b-mono), UniXCoder (Guo et al., 2022a) with 125M parameters (microsoft/unixcoder-base), and StarCoder+ (Li et al., 2023b) with 15.5B parameters (bigcode/starcoderplus). The reason for selecting these models is that they cover a variety of parameter sizes, model architectures, and pre-training processes. We leave all parameters of the models at their defaults, except for the maximum new tokens parameter, which we set to 256 to allow for longer completions. As a baseline, we query the models with a prompt that contains all the code preceding the cursor. In case this prompt exceeds the maximum prompt size of 2,048 tokens, we truncate the prompt from the beginning.
Test Generation
For the test generation task, we build upon TestPilot and use GPT-3.5-turbo-0125 for the LLM as the OpenAI GPT models are already integrated into TestPilot and require minimal effort to run. As a baseline, we use the original TestPilot implementation with a one-hour time limit per package, 130k token limit per package, and four completions per prompt with temperature 0.1. To make the comparison fair, we set the token limit of De-Hallucinator to the amount of tokens used by the baseline, and the number of completions to four with temperature 0.1. This prevents De-Hallucinator being advantaged with generating more tokens. Moreover, we cache the model outputs so that the same prompts return the same completion for De-Hallucinator and the baseline.
5.1.3. Datasets
Table 2. List of projects used for evaluation.
Project (owner/name) Commit LoC Stars
Python projects used for code completion
graphql-python/graphene 57cbef6 9,484 7.8k
geopy/geopy ef48a8c 10,000 4.3k
nvbn/thefuck ceeaeab 12,181 83.3k
aaugustin/websockets∗∗ ba1ed7a 14,186 5k
arrow-py/arrow 74a759b 14,402 8.6k
lektor/lektor be3c8cb 16,852 3.8k
Parsely/streamparse aabd9d0 26,214 1.5k
Supervisor/supervisor ca54549 29,860 8.3k
mwaskom/seaborn f9827a3 37,367 12.1k
psf/black ef6e079 106,005 37.6k
scikit-learn/scikit-learn f3c6fd6 193,863 58.5k
JavaScript projects used for test generation
node-red/node-red 29ed5b2 60 18.8k
winstonjs/winston c63a5ad 496 22.2k
prettier/prettier 7142cf3 916 48.5k
tj/commander.js 83c3f4e 1,134 26.3k
js-sdsl/js-sdsl 055866a 1,198 0.7k
goldfire/howler.js 003b917 1,319 23.1k
websockets/ws b73b118 1,546 21.2k
handlebars-lang/handlebars.js 8dc3d25 2,117 17.8k
petkaantonov/bluebird df70847 3,105 20.4k
hapijs/joi 5b96852 4,149 20.7k
Unitech/pm2 a092db2 5,048 40.9k
11ty/eleventy e71cb94 5,772 16.4k
As of June 5, 2024
∗∗ Has been moved to python-websockets/websockets
Code Completion